### v1

# Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger

Abstract
Large language models (LLMs) have shown remarkable emergent capabilities, transforming the execution of functional tasks by leveraging external tools for complex problems that require specialized processing or real-time data. While existing research expands LLMs access to diverse tools (e.g., program interpreters, search engines, weather/map apps), the necessity of using these tools is often overlooked, leading to indiscriminate tool invocation. This naive approach raises two key issues:(1) increased delays due to unnecessary tool calls, and (2) potential errors resulting from faulty interactions with external tools. In this paper, we introduce meta-cognition as a proxy for LLMs self-assessment of their capabilities, representing the model’s awareness of its own limitations. Based on this, we propose MeCo, an adaptive decision-making strategy for external tool use. MeCo quantifies metacognitive scores by capturing high-level cognitive signals in the representation space, guiding when to invoke tools. Notably, MeCo is fine-tuning-free and incurs minimal cost. Our experiments show that MeCo accurately detects LLMs’ internal cognitive signals and significantly improves tool-use decision-making across multiple base models and benchmarks.
Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger
Wenjun Li*, Dexun Li*, Kuicai Dong, Cong Zhang, Hao Zhang, Weiwen Liu, Yasheng Wang, Ruiming Tang, and Yong Liu Huawei Noah’s Ark Lab {li.wenjun, lidexun, zhang.hao3, liuweiwen8, wangyasheng, tangruiming, liu.yong6@huawei.com}
**footnotetext: Both authors contributed equally to this work.
1 Introduction
Equipping large language models (LLMs) with tool-use capabilities allows them to overcome their limitations by accessing external/real-time data (Komeili, 2021; Tang et al., 2023), domain-specific knowledge (He-Yueya et al., 2023; Schick et al., 2024), and advanced specialized functionalities (Yang et al., 2023; Gao et al., 2023; Lu et al., 2024), thereby enabling them to handle more complex tasks beyond their inherent capabilities. While prior research has focused on expanding tool diversity (Qin et al., 2023; Hao et al., 2024) and optimizing their use (Patil et al., 2023; Shen et al., 2024), the decision-making process for determining when tools are necessary remains underexplored.
Indiscriminate invocation of external tools leads to two major issues: (1) increased latency (Qu et al., 2024; Wang et al., 2024), as external tools, such as search engine, typically operates slower compared to relying on internal knowledge retrieval, and (2) robustness risks, where dependence on external APIs increases the likelihood of errors due to tools malfunction or unnecessary tool-use (Qin et al., 2023; Lu et al., 2024; Wu et al., 2024).
Figure 1: Overview of MeCo: Learned Meta-Cognition determines the necessity for tool use or retrieval by using a trained meta-cognition probe to detect the internal state of an LLM.
To address these limitations, we propose an adaptive tool-use strategy that improves decision-making in LLMs. Our approach, MeCo (Meta-Cognition-oriented trigger), enables LLMs to self-assess their capabilities and determine whether external tools are needed to address a user’s query. MeCo incorporates three key principles:
  • Meta-Cognition: We propose the concept of meta-cognition, which involves evaluating an LLM’s understanding of its capabilities based on its representation space. This self-awareness is crucial for minimizing unnecessary tool use.
  • Effective Strategy Utilization: We design an efficient strategy that leverages the quantified meta-cognition feedback to dynamically adjust tool use decisions, significantly improving decision accuracy compared to existing methods.
  • Generability: We empirically validate MeCo’s ability to generalize across a wide range of diverse scenarios. Additionally, we treat adaptive Retrieval-Augmented Generation (RAG) as a special case of tool use and demonstrate the superior effectiveness of MeCo in adaptive RAG compared to baseline methods.
Building on the Representation Engineering (RepE) framework (Zou et al., 2023), we develop a computationally efficient plug-in module to assess LLM meta-cognition. Our analysis reveals that meta-cognition produces strong signals, which can enhance interpretability and guide tool-use decisions. As shown in Figure 1, MeCo employs a dual-threshold policy to distinguish between strong and weak meta-cognitive signals, refining decision-making when uncertainty arises.
In summary, our contributions are four-fold: 1) We introduce the concept of adaptive tool use, which enhances both the efficiency and robustness of existing tool use paradigms in LLMs. 2) We integrate adaptive tool use and adaptive RAG within a unified framework, with their activation driven by a shared strategy based on meta-cognition detection. 3) We build a new benchmark, MeCa, to evaluate the effectiveness of our method. 4) We empirically demonstrate that MeCo significantly enhances the model’s awareness in tool utilization and RAG processes.
Figure 2: Pipeline for training the meta-cognition probe.
2 Background
Recent studies have explored LLMs’ internal representations to improve interpretability and understand their implicit beliefs (Bricken et al., 2023; Levinstein and Herrmann, 2024). Research (Zou et al., 2023; Liu et al., 2023a) shows that features like happiness, honesty, and confidence correspond to distinct, linearly separable directions in the representation space. Figure 2 illustrates the pipeline for training the corresponding feature probe. To detect these signals, contrastive instruction pairs can be used to induce their emergence.
Building on these insights, we can capture and control high-level functions $f$ (e.g., honesty) in model responses. We follow Zou et al. (2023) to design an experimental prompt $T^{+}_{f}$ that requires function execution and a reference prompt $T^{-}_{f}$ that does not. The instruction template is:
USER: $\langle$ instruction $\rangle$ $\langle$ experimental/reference prompt $\rangle$ ASSISTANT: $\langle$ output $\rangle$
For a function $f$ and model $M$, given instruction-response pairs $(q_{i},a_{i})$ in set $S$ and denoting a response truncated after token $k$ as $a^{k}_{i}$, we collect internal representations for the experimental and reference sets:
$A^{\pm}_{f}=\left\{\text{Rep}(M,T^{\pm}_{f}(q_{i},a^{k}_{i}))[-1]\text{ }|\text{ }(q_{i},a_{i})\in S\right\}$ (1)
where Rep represents the representation obtaining operation, $[-1]$ denotes the last token representation of $a^{k}$, and $A^{\pm}_{f}$ are the resulted activations consist of individual vectors.
Our goal is to learn a linear model to predict the direction of the function $A^{\pm}_{f}$ based on internal representations. Specifically, we apply PCA (Maćkiewicz and Ratajczak, 1993) in an unsupervised manner to pairwise difference vectors, deriving the first principal component $v_{f}$ (referred to as the probe) to identify function direction in the model’s responses. Equation 1 is applied at each layer of $M$ to derive layer-wise probes which are then used to interact with the LLM’s representations to monitor and control its behavior.
3 Approach
We define meta-cognition in LLMs as follows:
Definition 1
Meta-cognition refers to an LLM’s ability to assess and regulate its own knowledge and limitations, enabling informed decision-making about task execution, including when to rely on external tools or resources.
In the context of tool use, this involves assessing the model’s capabilities and limitations to determine whether a query can be answered independently or requires external tools, based on the query’s complexity and the sufficiency of the model’s internal knowledge.
To quantify meta-cognition, we train a probe that detects the model’s level of meta-cognitive awareness. This probe evaluates the rationale behind the model’s decision-making process, providing a score that reflects the model’s self-assessment accuracy. For instance, when the model receives a complex mathematical query, the meta-cognition probe assesses whether it correctly decides to solve the problem itself or delegate it to a calculator. A high meta-cognition score indicates accurate self-assessment, while a low score suggests either an incorrect attempt to solve it independently or unnecessary reliance on the tool.
3.1 Meta-Cognition Probe Extraction
Training a meta-cognition probe significantly differs from training probes for concepts like honesty or confidence. The latter concepts typically involve true-false statements about facts, such as "fire needs oxygen to burn" and "oxygen is harmful to human breathing". These statements are independent of user queries, meaning the model produces the same statements regardless of the user query.
In contrast, detecting the model’s internal cognition regarding tool use requires query-dependent responses. To achieve this, we employ leading proprietary LLM to generate user queries related to tool use and their corresponding responses (i.e., Yes/No responses with brief explanations). We then construct the training dataset following the procedures outlined in Section 2.
Notably, only a small dataset (of query-response pairs) suffices for strong probe performance. The analysis of the relationship between probe performance and the size of the training data is provided in Appendix D. Specifically, after collecting the instruction response pairs $(q_{i},a_{i})$, where $i$ denotes the index of the queries. We gather the sets of internal representations from the paired data and compute $A^{\pm}_{f}$ according to Eq. 1, and then apply PCA to the input $\{(-1)^{i}(A_{f,i}^{+}-A_{f,i}^{-})\}$ to obtain the first principal component $\nu_{f}$ as the meta-cognition probe. This principal component vector identifies the direction that represents the underlying meta-cognition concept.
After the training procedure described above, we obtain the meta-cognition probes across all model layers (e.g., 32 probes for Llama-3-8B). We compare our meta-cognition probe to existing proposed honesty Zou et al. (2023) and confidence probes Liu et al. (2024a) by evaluating the intermediate classification accuracy on held-out examples where the model is instructed to exhibit either honest/confident/strong meta-cognition or dishonest/lacking confidence/weak meta-cognition. The results of the probes are shown in Figure 3. As illustrated in Figure 3, our meta-cognition probe achieves near-optimal accuracy, significantly outperforming prior work.
Figure 3: Comparison between different probes. Note that -1 means the last layer in the LLMs.
Figure 4: Overview of benchmarks: Distribution of Metatool, MeCa-Tool, and MeCa-RAG categories. Metatool and MeCa-Tool assess the necessity of tool use, while MeCa-RAG evaluates the necessity of RAG interactions.
3.2 Decision-Making Strategy based on Meta-Cognition
With an accurate meta-cognition probe, we design a decision-making strategy based on detection results. For a given query, the LLM generates a response of $m$ tokens, each assigned a meta-cognition score across $n$ layers, forming a detection array of size $(m,n)$. The final decision—"Yes" (use external tools/RAG) or "No" (respond independently)—is derived from this array.
Reducing $m$ to 1.
We examine various prompting strategies (detailed in Appendix C) and find that the Yes/No+Explanation strategy, where the model answers with "Yes" or "No" followed by a brief explanation, yields the best performance. Therefore, we focus on the first token of the model’s response as it provides a clear signal of whether the model decides to rely on external tools. Extracting the meta-cognition score of the first token to represent the whole response simplifies our decision-making process, as calculating an overall meta-cognition score for the entire response is challenging due to varying response lengths and content across different queries. Since the model always responds with "Yes" or "No" as the first token, basing the trigger mechanism on the first token’s meta-cognition score is both reasonable and practical.
Reducing $n$ to 1.
In Zou et al. (2023) and Liu et al. (2024a), a mean score from multiple probes’ results is usually used to represent the token’s final quantification. However, our experiments show that scores predicted by different probes vary significantly, and simply averaging multiple scores does not yield accurate results. We found that probes in shallower layers (e.g., layer -5 to -2) tend to be more effective, with appropriate score distributions, ranges, and lower variances. Therefore, we select a single probe with the highest classification accuracy in the layer -5 to -2 (as shown in Figure 3) as the final predictor.
With meta-cognition scores distilled into a single scalar value, we apply the dual-thresholding strategy shown in Figure 1 to determine the optimal thresholds, $l_{yes}$ and $l_{no}$, using validation data. These thresholds are then applied to the test data.
4 Benchmark-MeCa
We evaluate MeCo using Metatool (Huang et al., 2023) and introduce a new benchmark, Meta-Cognitive Tool Assessment (MeCa), where each query undergoes human review. MeCa extends Metatool by incorporating a broader range of scenarios to assess tool usage and adaptive RAG.
Metatool comprises 1,040 queries designed to evaluate whether LLMs recognize when to use external tools. It includes 166 tools sourced from OpenAI’s plugin list (OpenAI, 2023). In Metatool, LLMs must decide on tool usage based solely on user queries, without tool names or descriptions. However, Metatool has limitations: 1) queries lack supplementary information or explicit tool provisions, whereas real-world tasks involve more complex intents and diverse requirements; 2) it primarily focuses on single-turn tool usage decisions. To address these gaps and ensure a more robust evaluation of MeCo, we developed MeCa, including two main components, MeCa-Tool and MeCa-RAG.
MeCa-Tool enhances Metatool by expanding tool-related assessments into three key categories:
  • Tool Usage Assessment: Evaluates whether an LLM should invoke external tools.
    • Queries solvable by the LLM without tools.
    • Queries requiring one or more tools due to insufficient internal capabilities.
  • Provided Tool Evaluation: Tests the LLM’s ability to determine tool usage when provided with a predefined set of tools.
    • Cases where external tools are unnecessary.
    • Cases where essential tools are available and should be used.
    • Cases where required tools are missing.
  • Multi-turn Interaction: Assesses tool usage decisions in multi-turn dialogues, requiring adaptation to evolving contexts.
    • Cases where external tools are unnecessary.
    • Cases where essential tools are available and should be used.
    • Cases where required tools are missing.
Specifically, we create six evaluation tasks in MeCa-Tool to systematically assess an LLM’s ability to make tool-related decisions across different scenarios. Tasks 1 and 4 evaluate whether an external tool is necessary to solve a query. Tasks 2 and 5 test the LLM’s ability to determine the relevance of a provided tool for a given query, including cases where the tool is irrelevant. Tasks 3 and 6 further extend this evaluation by presenting multiple tools (ranging from 2 to 5) and requiring the LLM to select the appropriate one. Notably, Tasks 1–3 focus on single-turn scenarios, while Tasks 4–6 incorporate multi-turn dialogues, assessing how well the LLM adapts its tool-use decisions in evolving conversational contexts. MeCa-Tool significantly expands Metatool by covering six tasks with 7,000 queries, providing a more diverse and comprehensive evaluation framework. The query composition of Metatool and MeCa-Tool is illustrated in Figure 4. Table 5 in Appendix A provides detailed statistics for each task.
Method Pre Fine-tuning Post Fine-tuning
w/ ctx w/o ctx w/ ctx w/o ctx
LM3-8B Naive 61.9 58.3 82.1 80.8
$P_{\text{Yes}}$ 63.5 62.7 81.7 80.8
MeCo 65.0 74.0 84.3 82.3
LM3-70B Naive 84.6 68.8 86.0 77.7
$P_{\text{Yes}}$ 84.8 73.7 86.2 77.1
MeCo 85.4 79.6 87.3 81.2
Mist-7B Naive 69.0 68.5 89.2 86.0
$P_{\text{Yes}}$ 71.2 73.1 89.2 85.0
MeCo 75.4 74.7 90.2 86.5
GPT-4-turbo 84.4 61.3 - -
Table 1: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on Metatool. Note that we are unable to calculate $P_{\text{Yes}}$ or detect the internal states of proprietary LLMs such as GPT-4-turbo.
Method Task 1 Task 2 Task 3 Task 4 Task 5 Task 6
w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx
LM3-8B Naive 70.0 65.0 62.3 80.3 54.3 78.7 66.0 50.0 70.5 54.0 60.5 53.5
$P_{Yes}$ 74.0 67.0 78.0 80.7 66.0 81.3 66.0 62.0 72.0 71.5 62.0 64.5
MeCo 79.0 72.0 80.1 81.3 73.3 79.5 69.0 69.0 74.0 78.5 63.5 67.0
Mist-7B Naive 54.0 63.0 42.3 55.7 55.7 67.3 60.5 70.0 73.5 76.0 73.0 62.5
$P_{Yes}$ 54.0 63.0 45.0 60.0 56.7 70.7 66.5 71.0 73.0 76.0 73.5 63.0
MeCo 58.0 67.0 66.7 66.0 74.8 78.3 69.0 78.5 76.2 80.0 74.0 65.5
Table 2: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-Tool before fine-tuning. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx for brevity.
Method Task 1 Task 2 Task 3 Task 4 Task 5 Task 6
w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx
LM3-8B-sft Naive 69.0 80.0 53.3 61.0 59.0 68.7 74.0 77.0 71.0 78.5 78.5 83.0
$P_{Yes}$ 70.0 78.0 58.3 68.7 57.7 70.0 75.0 77.0 80.5 84.0 81.5 82.0
MeCo 69.0 80.0 59.9 70.3 60.0 73.4 75.0 84.5 79.5 82.0 80.0 86.5
Mist-7B-sft Naive 68.0 64.0 52.3 53.0 58.3 73.7 92.5 77.5 87.5 82.0 85.0 70.5
$P_{Yes}$ 69.0 63.0 55.3 62.3 61.0 75.0 92.5 80.5 87.5 83.0 86.5 78.0
MeCo 71.0 66.0 60.7 66.3 65.7 82.0 95.0 87.0 88.0 82.0 88.0 80.5
Table 3: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-Tool after fine-tuning on the crafted SFT data. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx for brevity.
Beyond tool use, MeCa-RAG also evaluates adaptive RAG—determining whether a query can be answered directly by the LLM or requires external retrieval. RAG can be seen as a special case of tool use, where the LLM’s internal knowledge is insufficient and necessitates using a search engine to access external data. MeCa-RAG consists of:
  • Positive RAG: Cases where retrieval is essential for answering complex queries or those requiring up-to-date information.
  • Negative RAG: Cases where LLMs can answer using its internal knowledge without retrieval.
To curate the MeCa-Tool dataset, we employ a meticulous approach that began with collecting diverse scenarios from various online corpora, ensuring our synthetic APIs and conversations were grounded in realistic contexts. Based on these scenarios, we designed 500 distinct synthetic APIs emulating real-world applications across various domains and generated queries by randomly sampling from this API pool—queries that either require tool invocation, rely on the LLM’s internal knowledge, or expose cases where the provided APIs are insufficient—followed by rigorous human verification for accuracy. For the MeCa-RAG dataset, we selected common fact-based data from the RepE (Zou et al., 2023) dataset (e.g., “The Earth orbits the Sun”) to generate negative RAG queries that do not require retrieval, while positive RAG queries were generated from crawled recent news unseen by LLMs, necessitating retrieval of new information.
By incorporating both tool usage and adaptive RAG, MeCa serves as a robust benchmark for assessing LLM decision-making in complex scenarios. Detailed benchmark statistics and creation methodology for MeCa-RAG can be found in Appendix A.
5 Experiment Setup
Baselines: We evaluate MeCo against two baselines: Naive and $P_{\text{Yes}}$. The Naive baseline determines "Yes" or "No" based solely on the first token generated by the LLM, where "Yes" represents a positive indication, i.e., requiring external tools, and vice versa. The $P_{\text{Yes}}$ baseline refines this approach by computing a Yes-score,
[…]
which ranges from 0 (full "No") to 1 (full "Yes"), with values near 0.5 indicating uncertainty. Instead of relying solely on the first token, $P_{\text{Yes}}$ learns an optimal threshold: scores above this threshold are classified as "Yes," while those below are classified as "No." Further details are in Section C.2.
Backbone LLMs: We employ Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and Llama-3-70B-Instruct as backbone models to evaluate MeCo. For brevity, we refer to them as LM3-8B, Mist-7B, and LM3-70B. Additionally, we fine-tune these models on a dataset of 4,000 tool-use query-response pairs generated by GPT-4-turbo, denoting the fine-tuned versions as LM3-8B-sft, Mist-7B-sft, and LM3-70B-sft.
Evaluation: The primary evaluation metric is decision accuracy—whether the model correctly identifies when external tools/RAG are genuinely needed. Additional metrics (precision, recall, etc.) are analyzed in Appendix C.
Prompting Strategies: We experimented with various prompting strategies, including "Yes/No" with or without explanations and Chain-of-Thought (CoT; Wei et al. (2022)). The best performance was achieved using the "Yes/No + Explanation" strategy, which is used throughout this paper. Detailed results for different prompting strategies are available in Appendix C.
Moreover, we employ two types of prompts in our experiments: 1) prompts with context, which include specific reasons why LLMs may require external tools to complete user tasks, plus five randomly sampled examples to assist the model in making decisions; and 2) prompts without context, a concise version containing only the instruction and query. Exact prompts and settings are detailed in Appendix D.
6 Experiments
We conduct extensive experiments to empirically reveal the effectiveness of MeCo on two benchmarks: Metatool and MeCa. Specifically, we evaluate MeCo in adaptive tool use on both Metatool and MeCa and in adaptive RAG on MeCa.
6.1 MeCo in Adaptive Tool Use
First, we present the distribution of meta-cognition scores collected from the pre- and post-fine-tuning models in Figure LABEL:fig:metacog_distribution. We compare the meta-cognition scores for correct and incorrect responses and visualize how these scores differentiate between correct and incorrect Yes/No answers. Our key observations and interpretations are as follows:
  • Clear Gap in Meta-Cognition Scores: In both pre- and post-fine-tuning experiments, there is a noticeable gap between the meta-cognition scores of correct and incorrect responses. Our decision-making strategy can identify and leverage this gap to distinguish between correct and incorrect Yes/No answers.
  • Higher Scores for Correct Yes, Lower for Correct No: Correct Yes responses generally have higher meta-cognition scores than incorrect Yes responses, while correct No responses have lower scores than incorrect No responses. This occurs because the meta-cognition score for Yes/No tokens depends on the token embedding. Therefore, the meta-cognition scores of different tokens are not directly comparable; the score of Yes should only be compared to other Yes scores, and the score of No should only be compared to other No scores.
In our experiment, we sampled a subset of queries from the Metatool benchmark to create training data for determining the optimal thresholds for $P_{\text{Yes}}$ and MeCo. We then applied these thresholds to the test queries in both Metatool and MeCa-Tool (Task1 and Task4). Due to the fundamental difference between the queries in Metatool and those in Task 2, Task 3, Task 5, and Task 6 in MeCa-Tool, we randomly sample 100 queries from each of these categories to serve as hold-out testing data. We fit the thresholds for both $P_{\text{Yes}}$ and MeCo using the remaining data. The complete evaluation results are summarized in Table 1, Table 2, and Table 3. We make two key observations:
1. Superiority of MeCo: On both benchmarks, MeCo significantly enhances the model’s naive decision accuracy regarding tool use, outperforming $P_{\text{Yes}}$ by a considerable margin, indicating the effectiveness of the meta-cognition-based trigger mechanism. Notably, MeCo’s superiority is consistent across multiple backbone models and various evaluation settings, including both with and without context, as well as pre- and post-fine-tuning.
Importantly, the improvement achieved with MeCo incurs minimal costs, as it involves a fine-tuning-free and easy-to-integrate module. Note that fine-tuning and MeCo are two orthogonal approaches, and MeCo can provide additional benefits to fine-tuned models. Moreover, fine-tuned models do not transfer well to "out-of-distribution" testing scenarios. For instance, we observed performance degradation in the fine-tuned LM3-8B on Task 2 and Task 3 of MeCa-Tool. In contrast, the improvement brought by MeCo is consistent and robust across various testing scenarios.
MeCo’s superiority on MeCa is particularly promising and significant. MeCa contains more complex and realistic queries and user-assistant interactions, closely mimicking real-world scenarios. This underscores MeCo’s applicability to real-world LLMs, highlighting its potential for practical deployment and effectiveness in diverse and realistic scenarios.
2. Transferability: The results of Task1 and Task4 in Table 2 and Table 3 indicate that $P_{\text{Yes}}$ and MeCo, when fitted on one benchmark, can effectively transfer to other benchmarks. It’s worth noting that Metatool and MeCa feature different tool sources and styles of queries. We hypothesize that the model’s internal cognition is model-dependent, and once fitted on one benchmark, the decision strategy (i.e., the thresholds) can be transferred to other testing datasets. Although it is always better to align the decision strategy with real testing data, MeCo demonstrates satisfactory performance even when directly transferred, highlighting its robustness and adaptability.
Model Method Accuracy
LM3-8B Naive 63.0
$P_{\text{Yes}}$ 75.0
MeCo 76.0
Mist-7B Naive 84.0
$P_{\text{Yes}}$ 84.0
MeCo 86.0
GPT-3.5-Turbo - 86.0
GPT-4-Turbo - 84.0
Table 4: Performance Comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-RAG.
6.2 MeCo in Adaptive RAG
We further evaluate the effectiveness of MeCo in the adaptive RAG task, where the LLMs need to determine whether or not to retrieve external information to address the user query. Typically, no reasons or examples are provided to the LLMs in adaptive RAG, and we follow this setting by providing no context in the prompts. The results in Table 4 further validate the effectiveness of MeCo in the adaptive RAG task, demonstrating its robustness as a trigger mechanism across various applications. Note that GPT-4-turbo has more up-to-date information and thus does not perform RAG as often as GPT-3.5-turbo and results in a lower accuracy on our benchmark.
7 Conclusion
This paper introduces the concept of adaptive tool use to advance existing tool learning paradigms, which typically rely on external tools without discrimination to address user queries. We present MeCo, a computationally efficient plug-in module that assesses LLM’s meta-cognitive states, using a probe to detect relevant signals and inform more accurate tool-use decisions. To support evaluation, we introduce a new benchmark, MeCa, specifically designed to evaluate LLMs’ awareness of tool use as well as the necessity for retrieval. We empirically validate the effectiveness of MeCo using both the Metatool and MeCa, demonstrating significant improvements in the model’s decision-making accuracy regarding the necessity for tool use and retrieval. Our findings suggest that by integrating meta-cognition into the tool usage framework, we can enhance the operational efficiency and decision-making capabilities of LLMs across diverse contexts.
8 Limitations
We have omitted a fine-grained evaluation in the MeCa benchmark that measures the model’s end-to-end performance in correctly using tools with appropriate parameters. This includes determining whether tool use is necessary, parameter filling, and answer generation, etc. These tasks require capabilities beyond tool-use decision-making and would involve significant human effort to evaluate the final outputs. Our work mainly focuses on introducing adaptive tool use and leveraging meta-cognition for self-assessment in deciding when tool use is necessary. Although applying MeCo to parameter filling for tool use in LLMs is feasible, it also introduces additional latency. How to better apply MeCo to improve the parameter filling accuracy and the end-to-end evaluation remains an important direction for future work.
Appendix A MeCa Statistics
A.1 MeCa Statistics
Table 5 summarizes the statistics of the MeCa. In Task1 and Task4, the positive queries require a specific external tool to address the user queries, while the negative queries require no external tools and can be solved by the LLM’s internal capabilities. In Task2 and Task5, we provide a tool name and its description along with the user query, asking the LLMs to determine whether they need to use this specific tool to address the user queries. The neutral queries in Task2 and Task3 indicate that these queries require external tools, but the provided tool is irrelevant to addressing the user query. In Task3 and Task6, we provide a list of tools (ranging from 2 to 5) along with the user query. For multi-turn queries, there is a dialogue between the user and the LLM assistant, where the assistant needs to determine whether it should rely on external tools to address the user query in the final round of the conversation.
Task Category Count
MeCa-Tool-Task1 Positive queries without tools 500
Negative queries without tools 500
MeCa-Tool-Task2 Positive queries with relevant tools 500
Negative queries with tools 500
Neutral queries with irrelevant tools 500
MeCa-Tool-Task3 Positive queries with a tool list 500
Negative queries with a tool list 500
Neutral queries with a tool list 500
MeCa-Tool-Task4 Multi-turn Negative queries without tools 500
Multi-turn Positive queries without tools 500
MeCa-Tool-Task5 Multi-turn Positive queries with relevant tools 500
Multi-turn Negative queries with tools 500
MeCa-Tool-Task6 Multi-turn Positive queries with a tool list 500
Multi-turn Negative queries with a tool list 500
MeCa-RAG Positive RAG 150
Negative RAG 150
Table 5: Tool Usage Categories and Counts
We directly transfer the $l_{yes}$ and $l_{no}$ thresholds of MeCo, fitted on the Metatool dataset, to Task1 and Task4 in MeCa-Tool, and present the results in Table 2 and Table 3. Because the rest of the tasks in MeCa-Tool are very different and more complex than the user queries in MetaTool, we randomly sample 100 queries from each category in Task2, Task3, Task5, and Task6, and use these queries as the hold-out testing data. We use the remaining data to fit the thresholds for $P_{yes}$ and MeCo.
A.2 MeCa Creation
MeCa-Tool
To curate the MeCa-Tool dataset, we employed a meticulous and structured approach that ensures the queries are relevant to current LLM capabilities. The process unfolded as follows:
  1. 1.
    Collection of diverse scenarios: We began by gathering a broad spectrum of domains and conversational scenarios from various online corpus. This initial step ensures that the subsequent generated synthetic APIs and conversations are grounded in realistic and diverse settings.
  2. 2.
    Synthetic APIs design: Leveraging the collected scenarios, we then synthetically design 500 distinct APIs by emulating examples found in real-world applications, ensuring that they span multiple domains.
  3. 3.
    Query generation: For each query, APIs are randomly sampled from our synthetic APIs pool. User queries are then constructed based on sampled APIs, which may: (i) Require the invocation of the provided APIs; (ii) Not require any tool invocation, relying solely on the LLM’s internal knowledge; or (iii) Involve cases where the provided APIs does not include the necessary tools to answer the query directly.
  4. 4.
    Human Verification: After the queries were constructed, they underwent a rigorous human review process. This critical step verified the validity and correctness of the data, ensuring that each query aligns with its intended category and meets quality standards.
MeCa-RAG
The dataset was constructed as follows: we selected a subset of fact-based data from the RepE dataset (Zou et al., 2023), which consists of common, well-known facts, such as “The Earth orbits the Sun.” These facts were used as model responses, and the leading proprietary LLM (i.e., GPT-4-turbo) was instructed to generate corresponding user queries. Since these queries involve common knowledge that is embedded in LLMs, they do not require retrieval and thus serve as negative RAG examples. For positive RAG examples, we scraped recent news articles from the past few months, ensuring that this content has not been seen by LLMs. We then followed a similar process as mentioned above to generate user queries based on the latest information. This process resulted in queries that require retrieval as they involve knowledge that is unknown or not yet integrated into the LLM’s training data. The detailed distribution of MeCa can be found in Figure 4.
Appendix B Related Work
Tool Use in LLMs
LLMs have progressed from understanding and generating human-like text to utilizing external tools based on natural language instructions. This evolution expands their application beyond basic conversational tasks to enable dynamic interactions across diverse functional domains, such as facility management and professional services (Patil et al., 2023; Liu et al., 2023b; Qin et al., 2023; Chen et al., 2023). For example, Toolformer (Schick et al., 2024) enables LLMs to use external tools via simple APIs through a supervised fine-tuning (SFT) model. Liu et al. (2024c) demonstrate strong executable functional API calls across different domains. ToolACE (Liu et al., 2024b) trained on synthesized data, achieves state-of-the-art results on the Berkeley Function-Calling Leaderboard (Yan et al., 2024), even with a relatively small model size of 8B parameters. Despite their growing popularity and capabilities, tool use in LLMs often depends on strategies like verbal feedback, which are hampered by the quality of the datasets used for fine-tuning. Several benchmarks/datasets have been developed to support tool use in a data-centric way, such as API-Bank (Li et al., 2023), which provides a set of tool-use dialogues with various APIs to assess the LLM’s tool use capabilities, Toolalpaca (Tang et al., 2023) constructs a comprehensive tool-use corpus derived from collected real-world APIs, designed specifically to fine-tune LLMs for better tool utilization. ToolBench (Qin et al., 2023) focuses on creating a synthetic instruction-tuning dataset for tool use. However, these methods rely solely on superficial textual information, without probing deeper into the LLM’s internal states to explain or justify when and why a tool should be called, resulting in an inability to accurately determine the optimal timing for tool invocation.
Adaptive RAG
RAG has shown success in supporting AI systems that require up-to-date information or access domain-specific knowledge, particularly where the scope of queries is not seen in the training data of LLMs (Lewis et al., 2020; Ren et al., 2023; Vu et al., 2023; Izacard et al., 2023). This paper is also consistent with the trend of towards adaptive RAG paradigm, which is designed to assess whether a query can be directly answered by the LLMs or requires external data retrieval (Asai et al., 2023; Jiang et al., 2023). Specifically, a simple query within the LLM’s knowledge should be directly answered by the LLMs themselves. On the other hand, for complex queries or questions about data they have not been trained on, RAG intervenes to prevent incorrect out-of-date answers or hallucination (Ji et al., 2023). This mechanism allows RAG to dynamically adjust operational strategies of retrieval-augmented LLMs by assessing the boundary of LLM’s self-knowledge and the complexity of the query, thereby minimizing unnecessary computational overhead when the queries are answerable by LLMs themselves. Similar to the LLMs function-calling, the decision of retrieval timing typically hinges on three primary methods: (i) explicit verbal feedback from LLMs (Ding et al., 2024), (ii) enhancements through fine-tuning (Asai et al., 2023), or (iii) probability-based metrics (Kadavath et al., 2022; Jiang et al., 2023). Specifically, He et al. (2021) proposed enhancing the retrieval time efficiency by computing the probability of the next token via interpolating an LLM with a distribution calculated from the $k$ nearest context-token pairs. Drozdov et al. (2022) further extend $k$NN-LM to the adaptive paradigm by assigning the interpolation coefficient according to the retrieval quality measured by semantic similarity. Asai et al. (2023) introduce Self-RAG to improve generation quality and factuality by enabling adaptive retrieval and self-reflection. In contrast, this paper conceptualizes RAG as an external tool and highlights the importance of understanding the internal states of an LLM when developing the retrieval policy.
Explainability of LLMs
However, there is a considerable discrepancy between LLM’s decision mechanisms (often based on verbalized responses) and their internal cognition (Zou et al., 2023). The internal workings of LLMs are usually unclear, and this lack of transparency poses unwanted risks in downstream decision-making. Therefore, understanding and interpreting LLMs is crucial for elucidating their behaviors and limitations. To address this challenge, various explanations that provide insights into the inner workings of LLMs have been proposed (Zhao et al., 2024): (i) Probing-based explanations: Probing uses vector representations to measure embedded knowledge (Peters et al., 2018; Jawahar et al., 2019) or examines specific knowledge during the LLM’s generation process (Li et al., 2022), (ii) Neuron-level explanation: neuron analysis identifies critical neurons that are essential for model’s performance (Antverg and Belinkov, 2021; Bills et al., 2023), (iii) representation engineering (RepE): RepE leverages techniques inspired by cognitive neuroscience to identify and enhance the transparency of LLMs by uncovering their internal cognitive states (Zou et al., 2023). In this paper, we aim to detect the internal cognition of LLMs, and intervene LLM’s decisions, i.e., ensuring more precise decisions on tool use and retrieval timing.
Appendix C Extended Results
C.1 Prompting Strategies
To determine the best prompting strategy for tool use, we explore five prompting strategies with multiple base models. The results are summarized in Table 6.
  1. 1.
    Yes/No + Explanation: The model first answers with "Yes" or "No" and then provides a brief explanation for its decision.
  2. 2.
    Yes/No: The model answers solely with "Yes" or "No," without providing any explanation.
  3. 3.
    No/Yes + Explanation: The model first answers with "No" or "Yes" and then provides a brief explanation for its decision.
  4. 4.
    No/Yes: The model answers solely with "No" or "Yes," without providing any explanation.
  5. 5.
    CoT (Chain of Thought): The model is instructed to think step-by-step, reasoning why it does or does not need external tools to address the user query, and finally concludes its decision with “Yes" or “No."
[…]
Note that there are no results for the CoT prompting strategy for the Mistral-7b-instruct-v0.3 model. Regardless of the prompts used, the model consistently responds with “Yes/No" at the beginning, followed by an explanation of its decision. This behavior effectively mirrors the Yes/No+Explanation prompting strategy. Based on Table 6, we make the following observations and provide corresponding analysis:
  1. 1.
    Yes/No + Explanation generally performs the best out of the five prompting strategies. This approach provides a clear decision followed by reasoning, enhancing the model’s reliability and user trust.
  2. 2.
    CoT is not performing as well as expected. Through close human examination, we found that CoT results in long, complex answers where the model might ultimately conclude with a decision that contrasts with its prior reasoning process. This phenomenon is referred to as reasoning inconsistency, a challenge also reported in the literature(Wei et al., 2022; Lyu et al., 2023). Specifically, LLMs sometimes generate the correct answer following an invalid reasoning path or produce a wrong answer after a correct reasoning process, leading to inconsistency between the derived answer and the reasoning process. In contrast, the "Yes/No-Explanation" prompting strategy does not suffer from this reasoning inconsistency in our experiments, thereby achieving better performance compared to CoT.
  3. 3.
    Yes/No prompting strategy works better than No/Yes prompting. We hypothesize this phenomenon is due to the data format in the pre-training data. For example, there are likely many more Yes/No answers and reasoning processes in the training data compared to No/Yes answers, influencing the model’s performance.
We adopt Llama-3-8b-instruct and Mistral-7b-instruct-v0.3 as our backbone models because they exhibit strong performance in adaptive tool use. We exclude Llama-2-7b-chat due to its poor performance and lack of discernment regarding the necessity of external tools. Additionally, we exclude Llama-3.1-8b-instruct as its performance is almost identical to that of Llama-3-8b-instruct.
Table 6: Performance Comparison of Different Prompting Strategies
Model Prompting Strategies Accuracy Precision Recall F1 Score
Llama-2-7b-chat Yes/No+Explanation 0.51 0.51 1.0 0.67
Yes/No 0.51 0.5 1.0 0.67
No/Yes+Explanation 0.52 0.51 1.0 0.67
No/Yes 0.51 0.5 1.0 0.67
CoT 0.51 0.5 0.99 0.67
Llama-3-8b-instruct Yes/No+Explanation 0.72 0.82 0.57 0.67
Yes/No 0.63 0.61 0.72 0.66
No/Yes+Explanation 0.52 0.51 0.99 0.67
No/Yes 0.5 0.5 1.0 0.67
CoT 0.62 0.59 0.84 0.69
Llama-3.1-8b-instruct Yes/No+Explanation 0.71 0.66 0.87 0.75
Yes/No 0.64 0.59 0.95 0.73
No/Yes+Explanation 0.57 0.54 0.97 0.69
No/Yes 0.53 0.51 0.99 0.68
CoT 0.63 0.62 0.91 0.71
Mistral-7b-instruct-v0.3 Yes/No+Explanation 0.74 0.68 0.89 0.77
Yes/No 0.70 0.64 0.92 0.75
No/Yes+Explanation 0.70 0.64 0.88 0.74
No/Yes 0.71 0.57 0.82 0.74
C.2 P(Yes) Approach
The $P_{\text{Yes}}$ baseline introduces a Yes-score, as defined in Section 5. This score provides a nuanced measure of the model’s confidence, refining the binary approach taken by the Naive baseline. The Yes-score spans from 00 to $1$, where a score of 00 signifies a definite "No" and a score of $1$ signifies a definite “Yes". Scores close to $0.5$ reflect lower certainty in the model’s response, signifying ambiguity in decision-making. By adjusting the model’s output in cases where the Yes-score is near $0.5$ to always “Yes/No" answer, we aim to enhance the accuracy of both tool use and RAG timing. We employ 2 to determine the optimal threshold $l$ for the Yes-score based on training data, which is then applied to the test data.
$\text{Decision}=\begin{cases}\text{Yes}&\text{if }\textit{Yes-score}>l\\ \text{No}&\text{if }\textit{Yes-score}\leq l\end{cases}$ (2)
C.3 Distribution of P(Yes) and Meta-Cognition Scores
Before delving into the analysis, we provide some background on the concept of calibration in the context of Large Language Models (LLMs). Calibration refers to the alignment between a model’s predicted probabilities and the actual likelihood of those predictions being correct. A well-calibrated model generates probability scores that accurately reflect the true probability of its predictions.
In Figure LABEL:fig:p_yes_distribution, we present the distribution of P(Yes) scores for both correct and incorrect Yes/No decisions. Our key observations are as follows:
  1. 1.
    When the model is given detailed instructions and few-shot examples, it demonstrates poor calibration. As illustrated in Figure LABEL:fig:p_yes_distribution(a), the distributions of P(Yes) scores for correct and incorrect decisions do not show a clear distinction.
  2. 2.
    Conversely, when the model lacks detailed context and must rely on its internal beliefs to make decisions, it exhibits improved calibration. In Figure LABEL:fig:p_yes_distribution(b), the peak of the distribution for correct scores clearly deviates from that of incorrect scores.
  3. 3.
    After fine-tuning, the model displays significantly better calibration, as shown in Figures LABEL:fig:p_yes_distribution(c) and (d). Most correct decisions have P(Yes) scores of either 1 (indicating "Yes") or 0 (indicating "No"), while the P(Yes) scores for incorrect decisions vary between 0 and 1.
C.4 Meta-Cognition Scores at Different Layers
We examine the meta-cognition scores at various layers in the model and visualize the results in Figure 14. We focus on the meta-cognition scores at layers -2, -5, -8, -11, and -15 because these layers exhibit the highest classification accuracy, where layer -1 refers to the last layer before the output. Notably, the meta-cognition scores at different layers have distinct value ranges and slightly different distributions. Therefore, it is not reasonable to simply average the scores from different layers as the final score for a token, which has been a common approach in other research works based on RepE. In this study, we use the meta-cognition score from the second-to-last layer as the final score, as this layer demonstrates the highest classification accuracy and effectively differentiates between correct and incorrect responses.
Figure 14: Distribution of meta-cognition scores for the first token at different layers. The results are collected using the Llama-3-8b model on the training data from the Metatool benchmark.
Appendix D Probe Training
D.1 Different Training Strategies
Although it increases the length of the instructions and thus may degrade the signal we are detecting, we found that it is much better to provide the model with the query in the instruction than solely instruct the model to follow the ground truth explanations. Therefore, we include the queries in the contrastive instructions below.
Figure 15: The classification accuracy of different probes trained with the query in the instruction and without the query in the instruction. Training data size is fixed as 2048 in this experiment.
D.2 Different Size of Training Data
We further examine how the size of the training data affects the outcomes of the meta-cognition probe. Specifically, we analyze the performance of the trained probes with varying sizes of training data, as illustrated in Figure 16 and Figure LABEL:fig:probe_n_train_rag. According to Equation 1, a sentence with 10 tokens can be used to create 10 training data pairs of experimental prompts and reference prompts. Typically, a brief explanation of why or why not to use external tools/RAG corresponds to around 30 to 50 tokens. Thus, a training data size of 256 requires fewer than 10 queries and their associated explanations.
Although different backbone models exhibit significantly varying classification accuracies—with Llama-3-8b achieving the highest and Llama-3-70b the lowest—we found that only a small amount of training data is sufficient to train a probe with near-optimal performance. We hypothesize that this is due to the linear nature of the PCA methods adopted in RepE.
Figure 16: Training data size vs classification accuracy of meta-cognition probe in adaptive tool use.
D.3 Contrastive Instructions for Various Probes
We used the following instruction pair to collect contrastive data and train the honesty probe. Specifically, we instructed the model to be both honest and untruthful when explaining its reasoning for the necessity of tool use tes
[…]
Similarly, we instructed the model to exhibit both confidence and unconfidence when we trained the confidence probe.
[…]
For the meta-cognition probe, we instruct the model to exhibit strong meta-cognition by being constantly aware of its own limitations and capabilities and accurately assessing whether an external tool is necessary. Conversely, with weak meta-cognition, the model is often unaware of its own limitations and capabilities and struggles to assess the necessity of tool use.
[…]
The meta-cognition instruction for Adaptive RAG is similar to that in the adaptive tool use setting, with the only difference being that we replace the necessity of tool use with the necessity of RAG.
[…]
Appendix E Prompts
E.1 Prompts in Adaptive Tool Use
We employ two types of prompts in our experiments: 1) prompts with context, which provide specific reasons for why LLMs may require external tools to complete user tasks. These prompts also include five randomly sampled examples to assist the model in making decisions; and 2) prompts without context, which are more concise and contain only the instruction and query. The exact prompts are provided below. Note that the example queries are randomly sampled in the Metatool benchmark and we follow their setup and don’t change the examples associated with queries.
Prompt with context. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. Typically, there are four reasons why you might need to use external tools: A. Solving issues with real-time or external data, databases, or APIs B. Handling specialized inputs/outputs C. Enhancing domain tasks beyond LLM’s capabilities D. User customization, personalization, and interaction If you think it’s necessary to use external tools, please respond with “Yes"; otherwise, respond with “No". Additionally, you should provide a very brief explanation for your answer. Here are some examples: Query: “Write an opinion piece about why diversity and inclusion is super important for the tech industry. The essay should be targeted at ’tech bros’, and should avoid alienating them, but instead appeal to their logic; it should explain how diversity and inclusion of women, immigrants, etc. could benefit them specifically." Answer: No Query: “Are there any loopholes that hackers can exploit on my website?" Answer: Yes Query: “Plan a weekly lunch menu for a school. Write down a main dish, a carbohydrate side dish, a vegetable side dish, and a dessert for each day." Answer: No Query: “Can you break down the main points of this TED talk for me? Here’s the YouTube link." Answer: Yes Query: “How’s the weather in London right now?" Answer: No
[…]
E.2 Prompts in Adaptive RAG
In adaptive RAG task, LLMs are typically not provided with any reasons or examples to help them make a decision. Following this setting, we conduct the experiments in adaptive RAG without providing context in the prompts as shown below.
[…]