### v2

# Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger

Abstract
Large language models (LLMs) have shown remarkable emergent capabilities, transforming the execution of functional tasks by leveraging external tools for complex problems that require specialized processing or up-to-date data. While existing research expands LLMs access to diverse tools (e.g., program interpreters, search engines, calculators), the necessity of using these tools is often overlooked, leading to indiscriminate tool invocation. This naive approach raises two key issues: increased latency due to unnecessary tool calls, and potential errors resulting from faulty interactions with external tools. In this paper, we introduce meta-cognition as a proxy for LLMs self-assessment of their capabilities, reflecting the model’s awareness of its own limitations. Based on this, we propose MeCo, an adaptive decision-making strategy for external tool use. MeCo quantifies metacognitive scores by capturing high-level cognitive signals in the representation space, guiding when to invoke tools. Notably, MeCo is fine-tuning-free and incurs minimal cost. Experiments across multiple backbone models and benchmarks show that MeCo reliably detects LLMs’ internal cognitive signals and significantly improves tool-use decision-making.
Adaptive Tool Use in Large Language Models with Meta-Cognition Trigger
[…]
1 Introduction
Equipping large language models (LLMs) with tool use capabilities allows them to overcome their limitations by accessing external or up-to-date information (Komeili, 2021; Tang et al., 2023), domain-specific knowledge (He-Yueya et al., 2023; Schick et al., 2024), and advanced specialized functionalities (Yang et al., 2023; Gao et al., 2023; Lu et al., 2024), thereby enabling them to handle more complex tasks beyond their inherent abilities. While prior research has focused on expanding the tool arrays (Qin et al., 2023; Hao et al., 2024) and optimizing their use (Patil et al., 2023; Shen et al., 2024), the decision-making process for determining when tools are necessary remains underexplored.
Naively invoking tools indiscriminately leads to two major issues: (1) increased latency (Qu et al., 2024; Wang et al., 2024), as external tools, such as search engine, typically operates slower compared to relying on internal knowledge retrieval, and (2) robustness risks, where dependence on external APIs increases the likelihood of errors due to tools malfunction or unnecessary tool use (Qin et al., 2023; Lu et al., 2024; Wu et al., 2024).
Figure 1: Overview of MeCo: Learned Meta-Cognition determines the necessity for tool use or retrieval by using a trained meta-cognition probe to detect the internal state of an LLM.
To address these limitations, we propose an adaptive tool-use strategy that improves decision-making in LLMs. Our approach, MeCo (Meta-Cognition-oriented trigger), enables LLMs to self-assess their capabilities and decide whether external tools are needed to answer a given query. MeCo incorporates three key principles:
  • Meta-Cognition: We define meta-cognition in the context of LLM tool useas the model’s ability to evaluate its own competence based on internal representations. This self-awareness is crucial for minimizing unnecessary tool use.
  • Effective Strategy Utilization: We design an efficient strategy that leverages the quantified meta-cognition signals to dynamically adjust tool use decisions, significantly improving decision accuracy compared to baseline methods.
  • Generability: We demonstrate that MeCo generalizes well across diverse tasks and domains. Additionally, we treat adaptive Retrieval-Augmented Generation (RAG) as a special case of tool use and show MeCo outperforms standard baselines.
Building on the Representation Engineering (RepE) (Zou et al., 2023), we develop a computationally efficient plug-in module to assess meta-cognition in LLMs. Our analysis reveals that meta-cognition provides strong and interpretable signals that can effectively guide tool use decisions. As illustrated in Figure 1, MeCo employs a dual-thresholding policy to distinguish between strong and weak meta-cognitive signals, refining decision-making when uncertainty arises.
In summary, our contributions are four-fold: 1) We introduce the concept of adaptive tool use, which enhances both the efficiency and robustness of LLM tool use paradigms. 2) We unify adaptive tool use and adaptive RAG under a shared framework, with their activation driven by meta-cognition-based decision-making. 3) We build a new benchmark, MeCa, to systematially evaluate the effectiveness of our approach. 4) We empirically demonstrate that MeCo significantly enhances model awareness in both tool utilization and RAG scenarios.
Figure 2: Pipeline for training the meta-cognition probe.
2 Background
Recent studies have explored the internal representations of LLMs to improve interpretability and understand their implicit beliefs (Bricken et al., 2023; Levinstein and Herrmann, 2024). Prior work (Zou et al., 2023; Liu et al., 2023a) shows that abstract features like happiness, honesty, and confidence correspond to distinct, linearly separable directions in the representation space. Figure 2 illustrates the pipeline for training these feature probes. To detect these signals, contrastive instruction pairs are typically used to induce their emergence.
Building on these findings, we can capture and manipulate high-level functions $f$ (e.g., honesty) in model responses. We follow Zou et al. (2023) to design an experimental prompt $T^{+}_{f}$ that requires function execution and a reference prompt $T^{-}_{f}$ that does not. The instruction template is as follows:
USER: $\langle$ instruction $\rangle$ $\langle$ experimental / reference prompt $\rangle$ ASSISTANT: $\langle$ output $\rangle$
For a function $f$ and model $M$, given instruction-response pairs $(q_{i},a_{i})$ in set $S$ and denoting a response truncated after token $k$ as $a^{k}_{i}$, we collect internal representations for the experimental and reference sets:
$A^{\pm}_{f}=\left\{\text{Rep}(M,T^{\pm}_{f}(q_{i},a^{k}_{i}))[-1]\text{ }|\text{ }(q_{i},a_{i})\in S\right\}$ (1)
where Rep denotes the representation obtaining operation, [1][-1] is the last token representation of $a^{k}$, and $A^{\pm}_{f}$ are the resulted activations consist of individual vectors.
The goal is to learn a linear model to predict the direction of the function $A^{\pm}_{f}$ based on internal representations. Specifically, we apply PCA (Maćkiewicz and Ratajczak, 1993) in an unsupervised manner to pairwise difference vectors, deriving the first principal component $v_{f}$ (referred to as the probe) to identify function directions in the representation space. Equation (1) is applied at each layer of $M$ to derive layer-wise probes which are then used to interact with the LLM’s representations to monitor and control its behavior.
3 Approach
We define meta-cognition in LLMs as follows:
Definition 1
Meta-cognition refers to an LLM’s ability to assess and regulate its own knowledge and limitations, enabling informed decision-making about task execution, including when to rely on external tools or resources.
In the context of tool use, this involves assessing the model’s capabilities and limitations to determine whether a query can be answered independently or requires external tools, based on the complexity of the query and the sufficiency of model’s internal knowledge.
To quantify meta-cognition, we train a probe that detects the model’s level of meta-cognitive awareness. This probe evaluates the rationale behind the model’s decision-making process, providing a score that reflects the model’s self-assessment accuracy. For instance, when the model receives a complex mathematical query, the meta-cognition probe assesses whether it correctly decides to solve the problem itself or delegate it to a calculator. A high meta-cognition score indicates accurate self-assessment, while a low score suggests uncertainty in the model’s decision.
3.1 Meta-Cognition Probe Extraction
Training a meta-cognition probe differs fundamentally from training honesty or confidence probes. The latter probes are typically trained on true/false factual statements, e.g., "fire needs oxygen to burn" and "oxygen is harmful to human breathing". These statements are independent of specific user queries, meaning the model produces consistent internal representations regardless of the prompt context.
In contrast, detecting internal cognitive signals related to tool use requires query-dependent responses. To achieve this, we use leading proprietary LLMs to generate user queries related to tool use and their corresponding responses (i.e., binary Yes/No answers with brief explanations). We then construct the training dataset following the procedures outlined in Section 2.
Notably, only a small dataset suffices for strong probe performance. Appendix D provides an analysis of the relationship between probe performance and training data size. Specifically, after collecting the instruction-response pairs $(q_{i},a_{i})$, we extract the model’s internal representations and compute the contrastive representation differences $A^{\pm}_{f}$ according to Equation (1). We then apply PCA to the set of difference vectors $\{(-1)^{i}(A_{f,i}^{+}-A_{f,i}^{-})\}$ to obtain the first principal component $\nu_{f}$ as the meta-cognition probe, which identifies the direction of the underlying meta-cognition signal.
After the above procedures, we obtain meta-cognition probes across all model layers (e.g., 32 probes for Llama-3-8B). We evaluate our meta-cognition probe against existing probes for honesty Zou et al. (2023) and confidence Liu et al. (2024a) by measuring the intermediate classification accuracy on held-out examples where the model is instructed to exhibit either honest/confident/strong meta-cognition or dishonest/lacking confidence/weak meta-cognition. Intermediate classification accuracy refers to the probe’s ability to correctly predict the directional alignment, i.e., whether the token representation points in a positive or negative direction along the concept axis (honesty/confidence/meta-cognition). This metric does not reflect final tool-use decision accuracy. The results shown in Figure 3 indicates our meta-cognition probe achieves near-optimal accuracy, significantly outperforming prior probes.
Figure 3: Comparison between different probes. Note that -1 means the last layer in the LLMs.
3.2 Decision-Making Strategy based on Meta-Cognition
With an accurate meta-cognition probe, we design a decision-making strategy based on the detection results. For a given query, the LLM generates a response of $m$ tokens, each with a meta-cognition score across $n$ layers, forming a detection array of size $(m,n)$. The final decision—"Yes" (use external tools/RAG) or "No" (respond independently)—is derived from this array.
Reducing $m$ to 1.
We examine various prompting strategies (detailed in Appendix C) and find that the Yes/No+Explanation strategy, where the model answers with "Yes/No" followed by a brief explanation, yields the best performance. Hence, we focus on the first token of the model’s response, as it provides a clear signal of whether the model decides to rely on external tools. Extracting the meta-cognition score of the first token to represent the whole response simplifies our decision-making process, since calculating an overall meta-cognition score for the entire response is challenging due to varying response lengths and content across different queries. As the model consistently responds with "Yes/No" as the first token, basing the trigger mechanism on the first token’s meta-cognition score is both reasonable and practical.
Reducing $n$ to 1.
In Zou et al. (2023) and Liu et al. (2024a), a mean score from multiple probes’ results is typically used to represent the token’s final quantification. However, our experiments show that scores predicted by different probes vary significantly, and simply averaging them does not yield accurate results. We found that probes in shallower layers (e.g., layers -5 to -2) tend to be more effective, with appropriate score distributions, ranges, and lower variances. Therefore, we select the single probe with the highest classification accuracy within layers -5 to -2 (as shown in Figure 3) as the final predictor.
Dual-Thresholding.
After distilling the meta-cognition scores into a single scalar value, we adopt a simple yet effective dual-thresholding strategy (as illustrated in Figure 1) to determine the optimal thresholds, $l_{yes}$ and $l_{no}$, using validation data. These thresholds are then directly applied to the test data. As shown in Figure LABEL:fig:metacog_distribution, we empirically observe that correct Yes responses tend to have higher meta-cognition scores than incorrect Yes responses, while correct No responses tend to have lower scores than incorrect No responses. Notably, in both pre- and post-fine-tuning experiments, there is a clear gap between the meta-cognition scores of correct and incorrect responses. Our decision-making strategy identifies and leverages this gap to make better decisions. This pattern arises because the meta-cognition score is influenced by the token embedding of Yes/No, making scores between the two classes not directly comparable. Consequently, Yes scores should only be compared with other Yes scores, and similarly for No scores.
Driven by this empirical observation, we retain the model’s decision when the meta-cognition score $\nu$ is higher than $l_{yes}$ for Yes decisions, or lower than $l_{no}$ for No decisions; and override the model’s decision when $\nu$ is lower than $l_{yes}$ for Yes decisions, or higher than $l_{no}$ for No decisions. We demonstrate the robustness and effectiveness of this dual-thresholding strategy through comprehensive evaluations on both the MetaTool and MeCa benchmarks in Section 6.
Figure 9: Overview of benchmarks: Distribution of MetaTool, MeCa-Tool, and MeCa-RAG categories. MetaTool and MeCa-Tool assess the necessity of tool use, while MeCa-RAG evaluates the necessity of RAG interactions.
4 Benchmark-MeCa
We evaluate MeCo using MetaTool (Huang et al., 2023) and introduce a new benchmark, Meta-Cognitive Tool Assessment (MeCa), in which each query undergoes human review. MeCa extends MetaTool by incorporating a broader range of scenarios to assess adaptive tool use and RAG.
MetaTool comprises 1,040 queries designed to evaluate whether LLMs recognize when to use external tools. In MetaTool, LLMs must decide on tool usage based solely on user queries, without tool names or descriptions. We identify the following limitations of MetaTool: 1) queries lack supplementary information or explicit tool provisions, whereas real-world tasks involve more complex intents and diverse requirements; 2) it primarily focuses on single-turn tool usage decisions.
To address these gaps and provide a more robust evaluation of MeCo, we introduce MeCa, which includes two main components: MeCa-Tool and MeCa-RAG. MeCa-Tool extends MetaTool by expanding tool-related assessments across three key categories:
  • Tool Use Assessment: Evaluates whether the LLM should invoke external tools.
    • Queries solvable by the LLM without tools.
    • Queries requiring one or more tools due to insufficient internal capabilities.
  • Provided Tool Evaluation: Tests the LLM’s ability to decide on tool usage when given a predefined set of tools.
    • Cases where external tools are unnecessary.
    • Cases where essential tools are available and should be used.
    • Cases where required tools are missing.
  • Multi-turn Interaction: Assesses tool use decisions in multi-turn dialogues that require adaptation to evolving contexts.
    • Cases where external tools are unnecessary.
    • Cases where essential tools are available and should be used.
    • Cases where required tools are missing.
Specifically, we create six evaluation tasks in MeCa-Tool to systematically assess an LLM’s ability to make tool-related decisions across various scenarios. Tasks 1 and 4 evaluate whether an external tool is necessary to solve a given query. Tasks 2 and 5 assess the LLM’s ability to determine the relevance of a provided tool, including cases where the tool is irrelevant. Tasks 3 and 6 further extend this evaluation by presenting multiple tools (ranging from 2 to 5) and requiring the LLM to select the appropriate one. Notably, Tasks 1–3 focus on single-turn settings, while Tasks 4–6 involve multi-turn dialogues, testing the LLM’s ability to adapt tool-use decisions to evolving conversational contexts. MeCa-Tool significantly expands MetaTool by covering six tasks with 7,000 queries, providing a more diverse and comprehensive evaluation framework. The query composition of MetaTool and MeCa-Tool is illustrated in Figure 9, and detailed task statistics are provided in Table 7 in Appendix A.
Beyond tool use, MeCa-RAG evaluates adaptive RAG—determining whether a query can be answered directly by the LLM or requires external retrieval. RAG is a special case of tool use, where the LLM’s internal knowledge is insufficient and necessitates using of a search engine to access external information. MeCa-RAG includes:
  • Positive RAG: Cases where retrieval is essential to answer complex queries or those requiring up-to-date information.
  • Negative RAG: Cases where the LLM can provide correct answers using its internal knowledge without retrieval.
To curate the MeCa-Tool dataset, we employed a meticulous approach that began with collecting diverse scenarios from various online corpora, ensuring our synthetic APIs and conversations were grounded in realistic contexts. Based on these scenarios, we designed 500 distinct synthetic APIs emulating real-world applications across various domains. We then generated queries by randomly sampling from this API pool—queries that either require tool invocation, rely on the LLM’s internal knowledge, or expose cases where the provided APIs are insufficient—followed by rigorous human verification for accuracy. For the MeCa-RAG dataset, we selected common fact-based data from the RepE (Zou et al., 2023) dataset (e.g., "The Earth orbits the Sun”) to generate negative queries that do not require retrieval, while positive queries were generated from crawled recent news unseen by LLMs, thereby necessitating retrieval of new information.
By incorporating both adaptive tool usage and adaptive RAG, MeCa serves as a robust benchmark for assessing LLM decision-making in complex scenarios. Detailed benchmark statistics and creation methodology for MeCa-RAG can be found in Appendix A.
5 Experiment Setup
Baselines: We evaluate MeCo against two baselines: Naive and $P_{\text{Yes}}$. The Naive baseline determines "Yes" or "No" based solely on the first token generated by the LLM, where "Yes" indicates the need for external tools, and "No" indicates otherwise. The $P_{\text{Yes}}$ baseline refines this by computing a Yes-score,
$\text{Yes-score}=\frac{P(\text{Yes }|\text{ Prompt})}{P(\text{Yes }|\text{ Prompt})+P(\text{No }|\text{ Prompt})}$
which ranges from 0 (full "No") to 1 (full "Yes"), with values near 0.5 indicating uncertainty. Instead of relying solely on the first token, $P_{\text{Yes}}$ learns an optimal threshold: scores above this threshold are classified as "Yes," while those below are classified as "No." Further details are in Section C.2.
Backbone LLMs: We employ Llama-3-8B-Instruct, Mistral-7B-Instruct-v0.3, and Llama-3-70B-Instruct as backbone models to evaluate MeCo. For brevity, we refer to them as LM3-8B, Mist-7B, and LM3-70B. Additionally, we fine-tune these models on a dataset of 4,000 tool-use query-response pairs generated by GPT-4-turbo, denoting the fine-tuned versions as LM3-8B-sft, Mist-7B-sft, and LM3-70B-sft. We include the 70B variant to demonstrate MeCo ’s effectiveness on large-scale models and its potential applicability in industry-grade deployments.
Evaluation: The primary evaluation metric is decision accuracy—whether the model correctly identifies when external tools/RAG are genuinely necessary. Additional metrics (precision, recall, etc.) are reported in Appendix C.
Prompting Strategies: We experimented with various prompting strategies, including "Yes/No" with and without explanations and Chain-of-Thought (CoT; Wei et al. (2022)). The best performance was achieved using the "Yes/No + Explanation" strategy, which is used throughout this paper. Full results for all prompting variants are provided in Appendix C. We further employ two types of prompts in our experiments: 1) prompts with context, which include specific reasons why LLMs require external tools to complete user tasks, plus five randomly sampled in-context examples to guide the model’s decisions; and 2) prompts without context, a concise version containing only the instruction and query. Exact prompt templates are detailed in Appendix D.
6 Experiments
We conduct extensive experiments to empirically reveal the effectiveness of MeCo on MetaTool and MeCa. Specifically, we assess MeCo in adaptive tool use on both MetaTool and MeCa-Tool and in adaptive RAG on MeCa-RAG.
6.1 Reduced Tool Invocations and Latency
Before delving into detailed benchmark performance, we first demonstrate that using MeCo in adaptive tool use and adaptive RAG reduces unnecessary tool invocations and overall latency. For this evaluation, we use the LM3-8B model.
We begin by evaluating the number of tool invocations on MetaTool, which contains an equal proportion of queries that require and do not require external tools. Table 1 summarizes the number of tool invocations made by each method. The Naive method invokes 296 tool calls before fine-tuning and 277 after fine-tuning. $P_{\text{Yes}}$ improves efficiency by reducing the number of invocations to 208 and 273, respectively. Notably, MeCo achieves the lowest number of tool invocations—206 before fine-tuning and 240 after—while maintaining the highest decision accuracy. These results highlight MeCo’s effectiveness in addressing the two key issues associated with indiscriminate tool use:
  • Reduced Latency: By minimizing unnecessary tool calls, MeCo lowers response latency, particularly for queries that do not require external tools (e.g., retrieval tools can take up to 14.8 seconds, as shown in Table 2).
  • Improved Robustness: Avoiding unnecessary tool invocations reduces reliance on potentially unreliable external APIs, thereby mitigating risks and malfunctions caused by unnecessary tool use.
Method Pre Fine-tuning Post Fine-tuning
Always Call 520 (50.0%) 520 (50.0%)
Naive 296 (61.9%) 277 (82.1%)
$P_{\text{Yes}}$ 208 (63.5%) 273 (81.7%)
MeCo 206 (65.0%) 240 (84.3%)
Table 1: Number of tool invocations and decision accuracy (in parentheses) of different methods on MetaTool.
Secondly, we evaluate the latency overhead introduced by MeCo on MeCa-RAG. This experiment adopts a standard RAG pipeline involving a retrieval tool, consisting of the following stages:
  • Pre-processing: Queries are rewritten to be clearer and more specific for search. If the query is complex, it is decomposed into simpler, related sub-queries.
  • Search Engine: Google Search is used to retrieve the top-5 most relevant web pages.
  • Post-processing: The model integrates the original query with the retrieved content to generate an answer. When needed, relevant details are summarized or extracted to directly address the user’s query.
Method Search Overhead (s) w/ ctx w/o ctx
Always Call 14.8 0.00 0.00
Naive 14.8 0.94 0.85
$P_{\text{Yes}}$ 14.8 0.98 0.89
MeCo 14.8 1.74 0.95
Table 2: Increased latency (in seconds) introduced by different methods when deciding whether to invoke a search on MeCa-RAG. With context and without context are abbreviated as w/ cts and w/o ctx for brevity.
The average response time for invoking the retrieval tool (including pre-processing, search, and post-processing) is 14.8 seconds. While MeCo introduces a slight latency overhead due to the computation of meta-cognition scores, this cost is minimal. As shown in Table 2, the additional latency incurred by MeCo is negligible relative to the overall retrieval time, particularly for searches without context. For instance, MeCo adds only 0.95 seconds (a 6.4% increase over $P_{\text{Yes}}$), which is marginal compared to the 14.8-second baseline. More importantly, MeCo avoids unnecessary tool invocations, particularly in multi-round retrieval settings, thereby improving overall response efficiency and mitigating the significant delays introduced by external tools.
Method Pre Fine-tuning Post Fine-tuning
w/ ctx w/o ctx w/ ctx w/o ctx
LM3-8B Naive 61.9 58.3 82.1 80.8
$P_{\text{Yes}}$ 63.5 62.7 81.7 80.8
MeCo 65.0 74.0 84.3 82.3
LM3-70B Naive 84.6 68.8 86.0 77.7
$P_{\text{Yes}}$ 84.8 73.7 86.2 77.1
MeCo 85.4 79.6 87.3 81.2
Mist-7B Naive 69.0 68.5 89.2 86.0
$P_{\text{Yes}}$ 71.2 73.1 89.2 85.0
MeCo 75.4 74.7 90.2 86.5
GPT-4-turbo 84.4 61.3 - -
Table 3: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$ and MeCo on the MetaTool benchmark. Note that we are unable to calculate $P_{\text{Yes}}$ or detect the internal states of proprietary LLMs such as GPT-4-turbo. As a result, GPT-4-turbo uses the Naive decision-making strategy.
Method Task 1 Task 2 Task 3 Task 4 Task 5 Task 6
w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx
LM3-8B Naive 70.0 65.0 62.3 80.3 54.3 78.7 66.0 50.0 70.5 54.0 60.5 53.5
$P_{\text{Yes}}$ 74.0 67.0 78.0 80.7 66.0 81.3 66.0 62.0 72.0 71.5 62.0 64.5
MeCo 79.0 72.0 80.1 81.3 73.3 79.5 69.0 69.0 74.0 78.5 63.5 67.0
Mist-7B Naive 54.0 63.0 42.3 55.7 55.7 67.3 60.5 70.0 73.5 76.0 73.0 62.5
$P_{\text{Yes}}$ 54.0 63.0 45.0 60.0 56.7 70.7 66.5 71.0 73.0 76.0 73.5 63.0
MeCo 58.0 67.0 66.7 66.0 74.8 78.3 69.0 78.5 76.2 80.0 74.0 65.5
Table 4: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$, and MeCo on MeCa-Tool before fine-tuning. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx, respectively.
Method Task 1 Task 2 Task 3 Task 4 Task 5 Task 6
w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx w/ ctx w/o ctx
LM3-8B-sft Naive 69.0 80.0 53.3 61.0 59.0 68.7 74.0 77.0 71.0 78.5 78.5 83.0
$P_{\text{Yes}}$ 70.0 78.0 58.3 68.7 57.7 70.0 75.0 77.0 80.5 84.0 81.5 82.0
MeCo 69.0 80.0 59.9 70.3 60.0 73.4 75.0 84.5 79.5 82.0 80.0 86.5
Mist-7B-sft Naive 68.0 64.0 52.3 53.0 58.3 73.7 92.5 77.5 87.5 82.0 85.0 70.5
$P_{\text{Yes}}$ 69.0 63.0 55.3 62.3 61.0 75.0 92.5 80.5 87.5 83.0 86.5 78.0
MeCo 71.0 66.0 60.7 66.3 65.7 82.0 95.0 87.0 88.0 82.0 88.0 80.5
Table 5: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-Tool after fine-tuning on the crafted SFT data. "With context" and "without context" are abbreviated as w/ ctx and w/o ctx, respectively.
6.2 MeCo in Adaptive Tool Use
In this experiment, we sampled a subset of queries from the MetaTool benchmark to construct validation data for determining the optimal thresholds for $P_{\text{Yes}}$ and MeCo. These thresholds were then applied to the test queries in both MetaTool and MeCa-Tool (Task 1 and 4). Given the fundamental differences between the queries in MetaTool and those in Task 2, 3, 5, and 6 of MeCa-Tool, we randomly sampled 100 queries from each of these categories to serve as held-out test data. The thresholds for both $P_{\text{Yes}}$ and MeCo were fit using the remaining data in each respective category. Comprehensive evaluation results are reported in Table 3, Table 4, and Table 5. We report the decision accuracy of GPT-4-turbo as a reference point to illustrate the difficulty level of the queries in the MetaTool benchmark. We highlight two key observations:
1. Superiority of MeCo: Across both benchmarks, MeCo significantly improves the model’s decision accuracy regarding tool use, outperforming the $P_{\text{Yes}}$ baseline by a notable margin. This highlights the effectiveness of the meta-cognition-based trigger mechanism. Crucially, MeCo’s advantages are consistent across different backbone models and evaluation settings—both with and without context, as well as before and after fine-tuning.
Importantly, the gains from MeCo come at minimal cost: it is a fine-tuning-free, easily integrable module. Note that fine-tuning and MeCo are two orthogonal approaches, and MeCo can provide additional benefits to fine-tuned models. Notably, fine-tuned models often struggle to generalize to "out-of-distribution" scenarios. For instance, we observed a performance drop in the fine-tuned LM3-8B on Tasks 2 and 3 of MeCa-Tool. In contrast, MeCo’s improvements are robust and consistently maintained across varied testing conditions.
The strong performance of MeCo on MeCa is particularly noteworthy, as MeCa presents more complex and realistic queries and user–assistant interactions that better reflect real-world usage. This underscores MeCo ’s potential for practical deployment and its effectiveness in diverse and realistic settings.
2. Transferability: Results from Tasks 1 and 4 in Table 4 and Table 5 show that both $P_{\text{Yes}}$ and MeCo, when fitted on one benchmark, can generalize effectively to others. Despite the differences in tool sources and query styles between MetaTool and MeCa benchmarks, both methods demonstrate promising cross-benchmark performance. We hypothesize that an LLM’s internal meta-cognition is primarily model-dependent, and once a decision strategy (i.e., thresholds) is learned on one dataset, it can generalize to others. Although aligning the decision strategy with the target data is always preferable, MeCo performs robustly even in direct transfer settings, highlighting its robustness and adaptability.
Model Method Accuracy (%)
LM3-8B Naive 63.0
$P_{\text{Yes}}$ 75.0
MeCo 76.0
Mist-7B Naive 84.0
$P_{\text{Yes}}$ 84.0
MeCo 86.0
GPT-4-Turbo Naive 84.0
Table 6: Decision accuracy (%) comparison between Naive, $P_{\text{Yes}}$ and MeCo on MeCa-RAG.
6.3 MeCo in Adaptive RAG
We further evaluate the effectiveness of MeCo in the adaptive RAG task, where LLMs must decide whether to retrieve external information to accurately answer a user query. In line with standard settings for adaptive RAG, no explicit reasons or examples are provided in the prompts—only the raw query is given. To contextualize the difficulty of the queries in MeCa-RAG, we report the decision accuracy of GPT-4-turbo as a reference. As shown in Table 6, MeCo consistently demonstrates robust performance across multiple models and outperforms the baseline methods, validating its utility as a general-purpose trigger mechanism beyond tool use.
7 Conclusion
We introduce the concept of adaptive tool use, which advances traditional tool-learning paradigms that often invoke external tools indiscriminately. We propose MeCo, a lightweight and fine-tuning-free plug-in module that leverages meta-cognitive signals within LLMs to make more informed decisions about tool invocation. MeCo uses a probing mechanism to assess internal representations, enabling more precise determinations of when external tools and retrieval are truly necessary. To support rigorous evaluation, we introduce MeCa, a new benchmark that captures both tool-use and retrieval-awareness across realistic and diverse scenarios. Through extensive experiments on both the MetaTool and MeCa benchmarks, we demonstrate that MeCo significantly enhances decision accuracy regarding the necessity of external tools and retrieval, across multiple models and settings. Our findings suggest that incorporating meta-cognition into the LLM tool-use framework can yield more efficient and reliable decision-making in practical applications.
8 Limitations
Our current evaluation of the MeCa benchmark primarily focuses on the decision-making component—whether tool use or retrieval is necessary—rather than the full end-to-end performance of LLMs in completing tasks using tools. This includes downstream tasks such as selecting the correct tool, filling in parameters appropriately, and generating accurate final responses. Evaluating these stages would require substantial human annotation effort and introduces complexity beyond the current scope. While MeCo could be extended to assist in downstream tasks such as parameter filling, doing so would increase inference latency. Exploring how to effectively integrate MeCo into the full tool-use pipeline—balancing accuracy, efficiency, and usability—remains an important avenue for future research.
Appendix A MeCa Statistics
A.1 MeCa Statistics
Table 7 summarizes the statistics of MeCa. In Tasks 1 and 4, positive queries require a specific external tool to address the user queries, while negative queries can be answered using the LLM’s internal capabilities alone. In Tasks 2 and 5, a tool name and its description are provided alongside the user query; the LLM must determine whether this specific tool is necessary to resolve the query. Neutral queries in Tasks 2 and 3 indicate that external tools are needed, but the given tool is irrelevant to solving the query. In Tasks 3 and 6, a list of tools (ranging from two to five) is provided with the user query. For multi-turn queries, a dialogue between the user and the assistant is presented, and the assistant must decide whether to rely on external tools to respond to the final turn.
Task Category Count
MeCa-Tool-Task1 Positive queries without tools 500
Negative queries without tools 500
MeCa-Tool-Task2 Positive queries with relevant tools 500
Negative queries with tools 500
Neutral queries with irrelevant tools 500
MeCa-Tool-Task3 Positive queries with a tool list 500
Negative queries with a tool list 500
Neutral queries with a tool list 500
MeCa-Tool-Task4 Multi-turn Negative queries without tools 500
Multi-turn Positive queries without tools 500
MeCa-Tool-Task5 Multi-turn Positive queries with relevant tools 500
Multi-turn Negative queries with tools 500
MeCa-Tool-Task6 Multi-turn Positive queries with a tool list 500
Multi-turn Negative queries with a tool list 500
MeCa-RAG Positive RAG 150
Negative RAG 150
Table 7: Tool use categories and counts.
We directly transfer the $l_{yes}$ and $l_{no}$ thresholds of MeCo, fitted on the MetaTool dataset, to Tasks 1 and 4 in MeCa-Tool, with results shown in Table 4 and Table 5. Since the remaining tasks in MeCa-Tool (Tasks 2, 3, 5, and 6) differ significantly from MetaTool and involve more complex user queries, we randomly sample 100 queries from each of these tasks as hold-out test data. The remaining data is used to fit the thresholds for $P_{\text{yes}}$ and MeCo.
A.2 MeCa Creation
MeCa-Tool
To construct the MeCa-Tool dataset, we followed a meticulous and structured process to ensure the queries are aligned with current LLM capabilities:
  1. 1.
    Collection of diverse scenarios: We began by gathering a wide range of domains and conversational scenarios from various online corpora. This step ensured that the subsequent synthetic APIs and conversations were grounded in realistic and diverse contexts.
  2. 2.
    Synthetic API design: Based on the collected scenarios, we synthetically created 500 distinct APIs, emulating patterns from real-world applications to span multiple domains.
  3. 3.
    Query generation: For each query, APIs were randomly sampled from the synthetic API pool. User queries were then constructed to reflect one of the following types: (i) Queries that require invoking the provided APIs; (ii) Queries solvable without external tools, relying solely on the LLM’s internal knowledge; or (iii) Queries where the provided APIs do not include the necessary tools to directly answer the query.
  4. 4.
    Human verification: All generated queries underwent a rigorous human review process to ensure validity, correctness, and proper categorization. This step guaranteed high-quality annotations and alignment with task definitions.
MeCa-RAG
The dataset was constructed as follows: we selected a subset of fact-based entries from the RepE dataset (Zou et al., 2023), which contains well-known facts such as "The Earth orbits the Sun." These facts were used as target responses, and the leading proprietary LLM (i.e., GPT-4-turbo) was prompted to generate corresponding user queries. Since these queries involve widely known information already embedded in LLMs, they do not require retrieval and thus serve as negative RAG examples. For positive RAG examples, we scraped recent news articles from the past few months, ensuring that the content was unlikely to have been seen during LLM training. Using a similar prompting process, we generated user queries based on this up-to-date information. These queries require retrieval, as they concern facts that are either unknown to or not yet encoded in the model’s training data. The overall distribution of MeCa is illustrated in Figure 9.
Appendix B Related Work
Tool Use in LLMs
LLMs have progressed from understanding and generating human-like text to utilizing external tools based on natural language instructions. This evolution expands their application beyond basic conversational tasks to enable dynamic interactions across diverse functional domains, such as facility management and professional services (Patil et al., 2023; Liu et al., 2023b; Qin et al., 2023; Chen et al., 2023; Zhang and Lee, 2025). For example, Toolformer (Schick et al., 2024) enables LLMs to use external tools via simple APIs through a supervised fine-tuning (SFT) model. Liu et al. (2024c) demonstrates strong executable functional API calls across different domains. ToolACE (Liu et al., 2024b) trained on synthesized data, achieves state-of-the-art results on the Berkeley Function-Calling Leaderboard (Yan et al., 2024), even with a relatively small model size of 8B parameters. Despite their growing popularity and capabilities, tool use in LLMs often depends on strategies like verbal feedback, which are hampered by the quality of the datasets used for fine-tuning. Several benchmarks/datasets have been developed to support tool use in a data-centric way, such as API-Bank (Li et al., 2023), which provides a set of tool-use dialogues with various APIs to assess the LLM’s tool use capabilities, Toolalpaca (Tang et al., 2023) constructs a comprehensive tool-use corpus derived from collected real-world APIs, designed specifically to fine-tune LLMs for better tool utilization. ToolBench (Qin et al., 2023) focuses on creating a synthetic instruction-tuning dataset for tool use. However, these methods rely solely on superficial textual information, without probing deeper into the LLM’s internal states to explain or justify when and why a tool should be called, resulting in an inability to accurately determine the optimal timing for tool invocation.
Adaptive RAG
RAG has shown success in supporting AI systems that require up-to-date information or access domain-specific knowledge, particularly where the scope of queries is not seen in the training data of LLMs (Lewis et al., 2020; Ren et al., 2023; Vu et al., 2023; Izacard et al., 2023; Dong et al., 2025). This paper is also consistent with the trend towards the adaptive RAG paradigm, which is designed to assess whether a query can be directly answered by the LLMs or requires external data retrieval (Asai et al., 2023; Jiang et al., 2023). Specifically, a simple query within the LLM’s knowledge should be directly answered by the LLMs themselves. On the other hand, for complex queries or questions about data they have not been trained on, RAG intervenes to prevent incorrect out-of-date answers or hallucination (Ji et al., 2023). This mechanism allows RAG to dynamically adjust operational strategies of retrieval-augmented LLMs by assessing the boundary of LLM’s self-knowledge and the complexity of the query, thereby minimizing unnecessary computational overhead when the queries are answerable by LLMs themselves. Similar to the LLMs function-calling, the decision of retrieval timing typically hinges on three primary methods: (i) explicit verbal feedback from LLMs (Ding et al., 2024), (ii) enhancements through fine-tuning (Asai et al., 2023), or (iii) probability-based metrics (Kadavath et al., 2022; Jiang et al., 2023). Specifically, He et al. (2021) proposed enhancing the retrieval time efficiency by computing the probability of the next token via interpolating an LLM with a distribution calculated from the $k$ nearest context-token pairs. Drozdov et al. (2022) further extends kkNN-LM to the adaptive paradigm by assigning the interpolation coefficient according to the retrieval quality measured by semantic similarity. Asai et al. (2023) introduces Self-RAG to improve generation quality and factuality by enabling adaptive retrieval and self-reflection. In contrast, this paper conceptualizes RAG as an external tool and highlights the importance of understanding the internal states of an LLM when developing the retrieval policy.
Explainability of LLMs
However, there is a considerable discrepancy between the LLM’s decision mechanisms (often based on verbalized responses) and their internal cognition (Zou et al., 2023). The internal workings of LLMs are usually unclear, and this lack of transparency poses unwanted risks in downstream decision-making. Therefore, understanding and interpreting LLMs is crucial for elucidating their behaviors and limitations. To address this challenge, various explanations that provide insights into the inner workings of LLMs have been proposed (Zhao et al., 2024): (i) Probing-based explanations: Probing uses vector representations to measure embedded knowledge (Peters et al., 2018; Jawahar et al., 2019) or examines specific knowledge during the LLM’s generation process (Li et al., 2022), (ii) Neuron-level explanation: neuron analysis identifies critical neurons that are essential for model’s performance (Antverg and Belinkov, 2021; Bills et al., 2023), (iii) representation engineering (RepE): RepE leverages techniques inspired by cognitive neuroscience to identify and enhance the transparency of LLMs by uncovering their internal cognitive states (Zou et al., 2023). In this paper, we aim to detect the internal cognition of LLMs, and intervene in LLM’s decisions, i.e., ensuring more precise decisions on tool use and retrieval timing.
Appendix C Extended Results
C.1 Prompting Strategies
To determine the best prompting strategy for tool use, we explore five prompting strategies with multiple base models. The results are summarized in Table 8.
  1. 1.
    Yes/No + Explanation: The model first answers with "Yes" or "No" and then provides a brief explanation for its decision.
  2. 2.
    Yes/No: The model answers solely with "Yes" or "No," without providing any explanation.
  3. 3.
    No/Yes + Explanation: The model first answers with "No" or "Yes" and then provides a brief explanation for its decision.
  4. 4.
    No/Yes: The model answers solely with "No" or "Yes," without providing any explanation.
  5. 5.
    CoT (Chain of Thought): The model is instructed to think step-by-step, reasoning why it does or does not need external tools to address the user query, and finally concludes its decision with "Yes" or "No."
Note that there are no results for the CoT prompting strategy for the Mistral-7b-instruct-v0.3 model. Regardless of the prompts used, the model consistently responds with "Yes/No" at the beginning, followed by an explanation of its decision. This behavior effectively mirrors the Yes/No+Explanation prompting strategy. Based on Table 8, we make the following observations and provide corresponding analysis:
  1. 1.
    Yes/No + Explanation generally performs the best out of the five prompting strategies. This approach provides a clear decision followed by reasoning, enhancing the model’s reliability and user trust.
  2. 2.
    CoT is not performing as well as expected. Through close human examination, we found that CoT results in long, complex answers where the model might ultimately conclude with a decision that contrasts with its prior reasoning process. This phenomenon is referred to as reasoning inconsistency, a challenge also reported in the literature (Wei et al., 2022; Lyu et al., 2023). Specifically, LLMs sometimes generate the correct answer following an invalid reasoning path or produce a wrong answer after a correct reasoning process, leading to inconsistency between the derived answer and the reasoning process. In contrast, the "Yes/No-Explanation" prompting strategy does not suffer from this reasoning inconsistency in our experiments, thereby achieving better performance compared to CoT.
  3. 3.
    Yes/No prompting strategy works better than No/Yes prompting. We hypothesize that this phenomenon is due to the data format in the pre-training data. For example, there are likely many more Yes/No answers and reasoning processes in the training data compared to No/Yes answers, influencing the model’s performance.
Chain of Thought Prompting. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. Typically, there are four reasons why you might need to use external tools: A. Solving issues with real-time or external data, databases, or APIs B. Handling specialized inputs/outputs C. Enhancing domain tasks beyond LLM’s capabilities D. User customization, personalization, and interaction Please think step by step, and provide a brief explanation for your decision at first. At last, please conclude with "Yes" if you need to use external tools, or "No" if you do not need external tools. {Few-shot Examples}
User query: {query}
Answer:
We adopt Llama-3-8b-instruct and Mistral-7b-instruct-v0.3 as our backbone models because they exhibit strong performance in adaptive tool use. We exclude Llama-2-7b-chat due to its poor performance and lack of discernment regarding the necessity of external tools. Additionally, we exclude Llama-3.1-8b-instruct as its performance is almost identical to that of Llama-3-8b-instruct.
Model Prompting Strategies Accuracy Precision Recall F1 Score
Llama-2-7b-chat Yes/No+Explanation 0.51 0.51 1.0 0.67
Yes/No 0.51 0.5 1.0 0.67
No/Yes+Explanation 0.52 0.51 1.0 0.67
No/Yes 0.51 0.5 1.0 0.67
CoT 0.51 0.5 0.99 0.67
Llama-3-8b-instruct Yes/No+Explanation 0.72 0.82 0.57 0.67
Yes/No 0.63 0.61 0.72 0.66
No/Yes+Explanation 0.52 0.51 0.99 0.67
No/Yes 0.5 0.5 1.0 0.67
CoT 0.62 0.59 0.84 0.69
Llama-3.1-8b-instruct Yes/No+Explanation 0.71 0.66 0.87 0.75
Yes/No 0.64 0.59 0.95 0.73
No/Yes+Explanation 0.57 0.54 0.97 0.69
No/Yes 0.53 0.51 0.99 0.68
CoT 0.63 0.62 0.91 0.71
Mistral-7b-instruct-v0.3 Yes/No+Explanation 0.74 0.68 0.89 0.77
Yes/No 0.70 0.64 0.92 0.75
No/Yes+Explanation 0.70 0.64 0.88 0.74
No/Yes 0.71 0.57 0.82 0.74
Table 8: Performance comparison of different prompting strategies.
C.2 P(Yes) Approach
The $P_{\text{Yes}}$ baseline introduces a Yes-score, as defined in Section 5. This score provides a nuanced measure of the model’s confidence, refining the binary approach taken by the Naive baseline. The Yes-score spans from 0 to $1$, where a score of 0 signifies a definite "No" and a score of $1$ signifies a definite "Yes". Scores close to $0.5$ reflect lower certainty in the model’s response, signifying ambiguity in decision-making. By adjusting the model’s output in cases where the Yes-score is near $0.5$ to always "Yes/No" answer, we aim to enhance the accuracy of both tool use and RAG timing. We employ Equation (2) to determine the optimal threshold $l$ for the Yes-score based on training data, which is then applied to the test data.
$\text{Decision}=\begin{cases}\text{Yes}&\text{if }\textit{Yes-score}>l\\ \text{No}&\text{if }\textit{Yes-score}\leq l\end{cases}$ (2)
C.3 Distribution of P(Yes) and Meta-Cognition Scores
Before delving into the analysis, we provide some background on the concept of calibration in the context of Large Language Models (LLMs). Calibration refers to the alignment between a model’s predicted probabilities and the actual likelihood of those predictions being correct. A well-calibrated model generates probability scores that accurately reflect the true probability of its predictions.
In Figure LABEL:fig:p_yes_distribution, we present the distribution of $P_{\text{Yes}}$ scores for both correct and incorrect Yes/No decisions. Our key observations are as follows:
  1. 1.
    When the model is given detailed instructions and few-shot examples, it demonstrates poor calibration. As illustrated in Figure LABEL:fig:p_yes_distribution(a), the distributions of $P_{\text{Yes}}$ scores for correct and incorrect decisions do not show a clear distinction.
  2. 2.
    Conversely, when the model lacks detailed context and must rely on its internal beliefs to make decisions, it exhibits improved calibration. In Figure LABEL:fig:p_yes_distribution(b), the peak of the distribution for correct scores clearly deviates from that of incorrect scores.
  3. 3.
    After fine-tuning, the model displays significantly better calibration, as shown in Figures LABEL:fig:p_yes_distribution(c) and (d). Most correct decisions have $P_{\text{Yes}}$ scores of either 1 (indicating "Yes") or 0 (indicating "No"), while the $P_{\text{Yes}}$ scores for incorrect decisions vary between 0 and 1.
C.4 Meta-Cognition Scores at Different Layers
We examine the meta-cognition scores at various layers in the model and visualize the results in Figure 13. We focus on the meta-cognition scores at layers -2, -5, -8, -11, and -15 because these layers exhibit the highest classification accuracy, where layer -1 refers to the last layer before the output. Notably, the meta-cognition scores at different layers have distinct value ranges and slightly different distributions. Therefore, it is not reasonable to simply average the scores from different layers as the final score for a token, which has been a common approach in other research works based on RepE. In this study, we use the meta-cognition score from the second-to-last layer as the final score, as this layer demonstrates the highest classification accuracy and effectively differentiates between correct and incorrect responses.
Figure 13: Distribution of meta-cognition scores for the first token at different layers. The results are collected using the Llama-3-8b model on the training data from the MetaTool benchmark.
Appendix D Probe Training
D.1 Different Training Strategies
Although it increases the length of the instructions and thus may degrade the signal we are detecting, we found that it is much better to provide the model with the query in the instruction than solely instruct the model to follow the ground truth explanations. Therefore, we include the queries in the contrastive instructions below.
Figure 14: The classification accuracy of different probes trained with the query in the instruction and without the query in the instruction. Training data size is fixed as 2048 in this experiment.
D.2 Different Size of Training Data
We further examine how the size of the training data affects the outcomes of the meta-cognition probe. Specifically, we analyze the performance of the trained probes with varying sizes of training data, as illustrated in Figure 15 and Figure LABEL:fig:probe_n_train_rag. According to Equation (1), a sentence with 10 tokens can be used to create 10 training data pairs of experimental prompts and reference prompts. Typically, a brief explanation of why or why not to use external tools/RAG corresponds to around 30 to 50 tokens. Thus, a training data size of 256 requires fewer than 10 queries and their associated explanations.
Although different backbone models exhibit significantly varying classification accuracies—with Llama-3-8b achieving the highest and Llama-3-70b the lowest—we found that only a small amount of training data is sufficient to train a probe with near-optimal performance. We hypothesize that this is due to the linear nature of the PCA methods adopted in RepE.
Figure 15: Training data size vs classification accuracy of meta-cognition probe in adaptive tool use.
D.3 Contrastive Instructions for Various Probes
We used the following instruction pair to collect contrastive data and train the honesty probe. Specifically, we instructed the model to be both honest and untruthful when explaining its reasoning for the necessity of tool use tes
Contrastive Instructions for training Honesty Probe Honest Instruction Imagine you are an honest intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Untruthful Instruction Imagine you are an untruthful intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Similarly, we instructed the model to exhibit both confidence and unconfidence when we trained the confidence probe.
Contrastive Instructions for training Confidence Probe Confident Instruction Imagine you are a confident intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Unconfident Instruction Imagine you are an unconfident intelligent assistant explaining why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
For the meta-cognition probe, we instruct the model to exhibit strong meta-cognition by being constantly aware of its own limitations and capabilities and accurately assessing whether an external tool is necessary. Conversely, with weak meta-cognition, the model is often unaware of its own limitations and capabilities and struggles to assess the necessity of tool use.
Contrastive Instructions for training Meta-Cognition Probe Strong Meta-Cognition Instruction in Adaptive Tool Use Imagine you are an intelligent assistant with strong meta-cognition, constantly aware of your own limitations and capabilities. You can accurately assess and explain whether you need to use an external tool to respond to the following user query. User Query: {query}
Answer:
Weak Meta-Cognition Instruction Imagine you are an assistant with weak meta-cognition, often unaware of your own limitations and capabilities. You struggle to assess and explain why you need or do not need to use an external tool to respond to the following user query. User Query: {query}
Answer:
The meta-cognition instruction for Adaptive RAG is similar to that in the adaptive tool use setting, with the only difference being that we replace the necessity of tool use with the necessity of RAG.
Contrastive Instructions for training Meta-Cognition Probe in Adaptive RAG Strong Meta-Cognition Instruction Imagine you are an intelligent assistant with strong meta-cognition, constantly aware of your own limitations and capabilities. You can accurately assess and explain whether you need to perform Retrieval Augmented Generation (RAG) to respond to the following user query. User Query: {query}
Answer:
Weak Meta-Cognition Instruction Imagine you are an assistant with weak meta-cognition, often unaware of your own limitations and capabilities. You struggle to assess and explain why you need or do not need to perform Retrieval Augmented Generation (RAG) to respond to the following user query. User Query: {query}
Answer:
Appendix E Prompts
E.1 Prompts in Adaptive Tool Use
We employ two types of prompts in our experiments: 1) prompts with context, which provide specific reasons for why LLMs may require external tools to complete user tasks. These prompts also include five randomly sampled examples to assist the model in making decisions; and 2) prompts without context, which are more concise and contain only the instruction and query. The exact prompts are provided below. Note that the example queries are randomly sampled in the MetaTool benchmark and we follow their setup and don’t change the examples associated with queries.
Prompt with context. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. Typically, there are four reasons why you might need to use external tools: A. Solving issues with real-time or external data, databases, or APIs B. Handling specialized inputs/outputs C. Enhancing domain tasks beyond LLM’s capabilities D. User customization, personalization, and interaction If you think it’s necessary to use external tools, please respond with "Yes"; otherwise, respond with "No". Additionally, you should provide a very brief explanation for your answer. Here are some examples: Query: "Write an opinion piece about why diversity and inclusion is super important for the tech industry. The essay should be targeted at ’tech bros’, and should avoid alienating them, but instead appeal to their logic; it should explain how diversity and inclusion of women, immigrants, etc. could benefit them specifically." Answer: No Query: "Are there any loopholes that hackers can exploit on my website?" Answer: Yes Query: "Plan a weekly lunch menu for a school. Write down a main dish, a carbohydrate side dish, a vegetable side dish, and a dessert for each day." Answer: No Query: "Can you break down the main points of this TED talk for me? Here’s the YouTube link." Answer: Yes Query: "How’s the weather in London right now?" Answer: No
User query: {query}
Answer:
Prompt without context. You are an intelligent agent, and you need to constantly be aware of your own limitations. I will provide you with a user’s query, and you should assess, based on your own capabilities, whether you need to use external tools to better address the user’s query. If you think it’s necessary to use external tools, please respond with "Yes"; otherwise, respond with "No". Additionally, you should provide a very brief explanation for your answer. User Query: {query}
Answer:
E.2 Prompts in Adaptive RAG
In adaptive RAG task, LLMs are typically not provided with any reasons or examples to help them make a decision. Following this setting, we conduct the experiments in adaptive RAG without providing context in the prompts as shown below.
Prompt without context. Imagine you are an intelligent assistant with strong meta-cognition, constantly aware of your own limitations and capabilities. Your task is to accurately assess and explain whether you need to perform Retrieval Augmented Generation (RAG) to respond to the following user query. If you determine that performing RAG is necessary, please respond with "Yes"; otherwise, respond with "No". Additionally, provide a very brief explanation for your decision. User Query: {query}
Answer: